# Leveraging high fluence and low pressure for pulsed laser deposition of high-mobility -Al2O3/SrTiO3 heterostructure growth

### v1

@@ Abstract
Four different mechanisms for oxygen-vacancy-driven 2DEG formation were discussed. These were SrTiO3-environment equilibrium, bombardment, reduction of SrTiO3 by -Al2O3, and oxygen recombination. Considering these mechanisms and comparing 40 samples grown with the four different protocols, pointed toward a low limit for high-mobility -Al2O3/SrTiO3 growth of the fluence at 3 J/cm2, and a necessity of using low-pressure to induce oxygen vacancies in the surface of SrTiO3 before deposition.
We thank R. T. Dahm, C. N. Lobato, V. Rosendal, and C. E. N. Petersen for helpful discussions.
T.H.O., A. C. and F.T. acknowledge support by research grant 37338 (SANSIT) from Villum Fonden. T.S.J. acknowledges support from the Novo Nordic Foundation Challenge Program, grant no. NNF21OC0066526 (BioMag).

### v2

@@ VII Conclusion
Four different mechanisms for oxygen-vacancy-driven 2DEG formation were discussed. These were SrTiO3-environment equilibrium, bombardment, reduction of SrTiO3 by -Al2O3, and oxygen recombination. -Al2O3 having a single-element cation allows high laser fluences during growth without affecting stoichiometry, providing an increased bombardment effect and available energy for the crystal growth. Considering the growth mechanisms and comparing 40 samples grown with the four different protocols, pointed toward a
need for low pressure to induce oxygen vacancies in the surface layers of SrTiO3 before deposition.
VIII Supplementary material
The supplementary material includes an overview of the 40 samples described in this investigation, linear fitting of the Hall resistance used to extract the carrier densities and mobilities in Fig. 3. Furthermore, it includes a detailed electronic transport analysis of one high-mobility sample (sample 1), including a 2-band and anomalous hall effect fitting.
IX Acknowledgements
We thank R. T. Dahm, C. N. Lobato, V. Rosendal, and C. E. N. Petersen for helpful discussions. T.H.O., A. C. and F.T. acknowledge support by research grant 37338 (SANSIT) from Villum Fonden. T.S.J. acknowledges support from the Novo Nordic Foundation Challenge Program, grant no. NNF21OC0066526 (BioMag).